# SoundMind: RL-Incentivized Logic Reasoning for Audio-Language Models

### v1

@@ Abstract
While large language models have shown reasoning capabilities, their application to the audio modality, particularly in large audio-language models (ALMs), remains significantly underdeveloped. Addressing this gap requires a systematic approach, involving a capable base model, high-quality reasoning-oriented audio data, and effective training algorithms. In this study, we present a comprehensive solution: we introduce the Audio Logical Reasoning (ALR) dataset, consisting of 6,446 text-audio annotated samples specifically designed for complex reasoning tasks. Building on this resource, we propose SoundMind, a rule-based reinforcement learning (RL) algorithm tailored to endow ALMs with deep bimodal reasoning abilities. By training Qwen2.5-Omni-7B on the ALR dataset using SoundMind, our approach achieves state-of-the-art performance in audio logical reasoning. This work highlights the impact of combining high-quality, reasoning-focused datasets with specialized RL techniques, advancing the frontier of auditory intelligence in language models. Our code and the proposed dataset are available at https://github.com/xid32/SoundMind.
@@
Xingjian Diao1, Chunhui Zhang1, Keyi Kong2, Weiyi Wu1, Chiyu Ma1, Zhongyu Ouyang1, Peijun Qing1, Soroush Vosoughi1, Jiang Gui1 1Dartmouth College, 2Shandong University xingjian.diao.gr@dartmouth.edu
@@ 1 Introduction
Figure 1: Overview of the end-to-end bimodal Natural Language Inference (NLI) task enabled by large audio-language models. The model receives input and produces output in either pure text or pure audio format.
@@
Large language models (LLMs) have made remarkable strides in reasoning capabilities through innovations such as Chain-of-Thought (CoT) prompting and specialized reasoning architectures. Recent models such as OpenAI’s o1 model (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025) have demonstrated exceptional performance on complex mathematical and programming tasks (Zhao et al., 2024; Team et al., 2025; Muennighoff et al., 2025). Particularly noteworthy is Deepseek-R1’s rule-based reinforcement learning approach, which enables emergent reasoning without relying on traditional frameworks like Monte Carlo Tree Search (Wan et al., 2024) or process reward models (Lightman et al., 2023). This general reasoning paradigm has successfully extended to the visual domain, where frameworks like Visual-CoT (Shao et al., 2024) have significantly enhanced multimodal models’ cognitive abilities for image and video reasoning.
@@
Despite these advances in text and visual reasoning, a significant gap exists in audio reasoning and generation capabilities. While audio-language models (ALMs) like Audio Flamingo (Kong et al., 2024) and Qwen2.5-Omni-7B (Xu et al., 2025) have made progress in audio understanding, end-to-end audio reasoning remains underdeveloped. This limitation stems primarily from two factors: (1) the simplicity of existing audio reasoning datasets (Suzgun et al., 2023; Kong et al., 2024), which often contain only brief textual labels without proper audio modality annotations, and (2) the technical challenges in maintaining reasoning coherence during long-duration audio generation. Current CoT methods applied to audio often lead to hallucinations and performance degradation when generating extended reasoning sequences. The lack of aligned audio-text annotations further impedes research on reasoning-driven audio generation, creating a bottleneck in developing ALMs with sophisticated reasoning capabilities.
@@
To address these challenges in audio reasoning (Figure 1), we introduce a comprehensive approach with two key components. First, we construct the Audio Logical Reasoning (ALR) dataset (Liu et al., 2023), containing 6,446 dual-modality samples with user content, CoT reasoning steps, final answers, and corresponding audio input/output. Derived from the Logi-QA 2.0-NLI dataset, ALR provides complete logical reasoning tasks with both content and response in text and audio modalities. Second, we propose the SoundMind algorithm, which builds upon the ALR dataset to address performance issues in long-form reasoning audio generation. Drawing inspiration from the Logic-RL framework (Xie et al., 2025a), SoundMind incorporates a strict format-based reward mechanism to prevent shortcut reasoning biased toward the textual modality. It leverages the REINFORCE++ algorithm (Hu, 2025) alongside the reward design principles from Deepseek-R1 for effective post-training. Our main contributions are as follows:
@@
  • ALR Dataset: We release a high-quality dual-modality dataset containing 6,446 samples with user content, CoT reasoning, answers, and corresponding audio input/output. The ALR dataset includes deep reasoning annotations in both text and audio formats, providing a valuable resource for RL-based training of audio-language models.
@@
  • SoundMind RL Algorithm: Taking into account the unique challenges of audio reasoning generation, we design a strict format-based reward mechanism that maintains reasoning coherence across modalities. By combining the improved REINFORCE++ algorithm with the reward design principles of Deepseek-R1, we fine-tune Qwen2.5-Omni-7B using reinforcement learning, achieving state-of-the-art performance across three types of modality input-output combinations on the ALR test set.
2 Related Work
Open-Source Datasets for Audio Reasoning.
Currently, there are very few publicly available datasets for audio reasoning. The audio datasets released in recent years typically use audio as a descriptive input, combined with text, and are fed jointly into ALMs. However, the final annotations in these datasets are purely textual (Suzgun et al., 2023; Kong et al., 2024; Ghosh et al., 2025). Recently, some researchers have begun to recognize the critical role of CoT prompting in audio reasoning datasets (Xie et al., 2025b). Nevertheless, our study is the first open-source dataset in which audio serves as the annotated modality for reasoning.
@@ 2 Related Work
ALR (Our) Speech 6446 1K
Table 1: Comparison of reasoning audio datasets. ALR is the only dataset providing both transcripts and chain-of-thought annotations in addition to audio, supporting step-by-step reasoning and multimodal evaluation.
@@
LLMs enhance their reasoning ability through in-context learning (ICL), which processes prompts within the surrounding context. CoT techniques further reinforce this capability. Prominent CoT approaches include Tree-of-Thought (ToT) (Yao et al., 2023), manually crafted few-shot CoT (Wei et al., 2022), and various automatic generation strategies (Jin et al., 2024). Recent works have also examined the necessity, theoretical underpinnings, and task-specific effectiveness of CoT reasoning (Sprague et al., 2025). OpenAI’s ol model (Jaech et al., 2024) has reignited interest in CoT prompting and is often paired with reinforcement learning-based training approaches (Hu, 2025; Xie et al., 2025a).
@@
Although current ALMs have made progress in comprehension and real-time response, their capability in CoT-style reasoning remains underexplored. Our study addresses this research gap by systematically investigating the application of CoT techniques within ALMs.
3 Audio Logic Reasoning Dataset
3.1 Main Features
We introduce the Audio Logic Reasoning (ALR) dataset, a novel benchmark specifically designed to advance audio-based reasoning research. Unlike most existing datasets that rely solely on textual annotations, ALR provides both audio-level and CoT annotations, enabling models to learn and reason directly from auditory inputs. The dataset focuses on speech-based scenarios, aligning with real-world applications such as spoken dialogue understanding and auditory commonsense reasoning.
As shown in Table 1, existing audio datasets such as CoTA (Xie et al., 2025b), AudioSkills (Ghosh et al., 2025), LongAudio (Ghosh et al., 2025), and Big Bench Audio (Suzgun et al., 2023) either lack audio-level annotations or do not provide CoT reasoning traces. For instance, while CoTA includes CoT-style annotations, it does not offer audio annotations/transcripts, limiting its ability to support models that operate directly on audio. Similarly, AudioSkills and LongAudio provide a large volume of audio samples but lack both reasoning annotations and step-by-step rationales.
In contrast, ALR is the only dataset in this comparison that includes both audio and CoT annotations. Although it contains fewer samples (6,446) than large-scale datasets such as AudioSkills, it is more focused on logic-oriented reasoning tasks and offers 1,074 hours of carefully annotated speech audio. This dual-modality annotation makes ALR a uniquely valuable resource for training and evaluating Logic-Aware ALMs in tasks requiring structured reasoning grounded in audio perception.
@@ 3.2 Data Generation Pipeline
Figure 2: Three-stage pipeline for constructing ALR samples. Step 1: Structured logical inputs are converted into natural prompts through colloquialization. Step 2: A large language model generates chain-of-thought reasoning and final answers. Step 3: Both input and output texts are synthesized into speech using a text-to-speech model, producing paired audio data aligned with logical reasoning.
The Audio Logic Reasoning (ALR) dataset is constructed via a structured pipeline that converts textual logical reasoning tasks into natural spoken audio interactions, enabling audio-based reasoning model development. The full process is in Figure 2.
We begin with logic problems from the LogiQA 2.0 NLI (Liu et al., 2023), which provides a structured triplet: a major premise, a minor premise, and a conclusion. These inputs are first processed through a colloquialization module to convert formal logic into natural, conversational language (e.g., Major premise is …, Let’s figure out the logical connection…). This step enhances the naturalness of the audio and simulates real user prompts.
The colloquialized content and instructions are combined into user content, which is then synthesized into speech audio using the high-quality text-to-speech model MegaTTS 3 (Jiang et al., 2025). This forms the input side of the dataset. The prompts we used can be found in Table 2.
Figure 3: An example in text from the ALR dataset. The sample consists of three components: (1) User Content, which includes a natural-language prompt and a structured logical triplet (major premise, minor premise, and conclusion); (2) Chain-of-Thought response generated by a large language model, which explains the logical reasoning behind the entailment decision; and (3) the final Answer. All components are available in both text and synthesized audio formats, supporting multimodal training and evaluation.
To generate the output, we use the large language model DeepSeek-R1 to produce CoT reasoning and final answers. These reasoning traces provide step-by-step explanations and improve supervision quality. The CoT outputs are also synthesized into audio using MegaTTS 3, producing the corresponding spoken response.
As a result, each ALR sample consists of a pair of spoken audio segments: one representing the user query and the other containing the reasoning and answer, as shown in Figure 3.
@@
Table 2: Instructional prompts at different positions during chain-of-thought generation. The system prompt defines the overall task, output format, and desired reasoning style for speech synthesis. Contextual prompts inserted before the major premise and after the conclusion further guide the model in interpreting logical relationships.
Training 2326 2858 182 1683 62.33 608.42
Test 296 360 158 1424 57.90 586.51
Validation 264 342 155 1426 57.85 586.39
Table 3: Detailed statistics of the Audio Logic Reasoning (ALR) dataset. “# Entailed” and “# Not-ent.” indicate the number of entailed and not-entailed samples, respectively. “Avg. Inp./Out. Tok.” denotes the average input/output tokens. “Avg. Inp./Out. Dur. (s)” indicates the average duration (in seconds) of input/output sequences.
4 SoundMind Algorithm
@@
To ensure the correctness of answer formatting, we first require that the token “Answer:” must appear within the last five characters of the model response. Considering the dual-modality annotation characteristics of the ALR dataset, we design two specific format scoring methods (denoted as $S_{\text{format}}$), calculated as follows:
@@
We additionally introduce a reward evaluation based on the reasoning length of the model response. This score ($S_{\text{len}}$) is computed by comparing the ratio of the model output length to the reference reasoning length. In light of the ALR dataset’s dual-modality annotations, we design two distinct length evaluation methods, defined as follows:
@@
$L$ denotes the length of text tokens, and $T$ denotes the length of audio tokens.
@@
In our setting, the policy $\pi_{\theta}$ corresponds to a large-scale ALM, which receives an audio question as input and produces a reasoning response in either text or audio form. To optimize the model with the above composite reward, we adopt the REINFORCE++ Hu (2025)—a clipped policy-gradient method that eliminates the need for a value (critic) network while leveraging PPO-style stability and sample efficiency. Specifically, REINFORCE++ updates the policy by maximizing the following objective:
@@ 5 Experiments
Figure 4: Histograms of audio durations across dataset splits and entailment labels. Subfigures (a–c) correspond to the training, test, and validation sets for “Entailed” samples; (d–f) show the same splits for “Not-entailed” samples.

### v2

@@ Abstract
While large language models have demonstrated impressive reasoning abilities, their extension to the audio modality, particularly within large audio-language models (LALMs), remains underexplored. Addressing this gap requires a systematic approach that involves a capable base model, high-quality reasoning-oriented audio data, and effective training algorithms. In this work, we present a comprehensive solution for audio logical reasoning (ALR) tasks: we introduce SoundMind, a dataset of 6,446 audio–text annotated samples specifically curated to support complex reasoning. Building on this resource, we propose SoundMind-RL, a rule-based reinforcement learning (RL) algorithm designed to equip audio-language models with robust audio–text reasoning capabilities. By fine-tuning Qwen2.5-Omni-7B on the proposed SoundMind dataset using SoundMind-RL, we achieve strong and consistent improvements over state-of-the-art baselines on the SoundMind benchmark. This work highlights the benefit of combining high-quality, reasoning-focused datasets with specialized RL techniques, and contributes to advancing auditory intelligence in language models. The code and dataset are publicly available at https://github.com/xid32/SoundMind.
@@
Xingjian Diao, Chunhui Zhang, Keyi Kong , Weiyi Wu, Chiyu Ma, Zhongyu Ouyang, Peijun Qing, Soroush Vosoughi, Jiang Gui Dartmouth College, Shandong University xingjian.diao.gr@dartmouth.edu [Uncaptioned image]SoundMind Dataset
@@ 1 Introduction
Figure 1: Overview of a SoundMind sample for the audio logical reasoning (ALR) task. Each instance contains User Content (top), including a natural-language prompt, structured logical triplets, and a question, as well as Response (bottom), which consists of step-by-step chain-of-thought reasoning and the final answer. Both components are provided in text and accompanied by the corresponding input–output speech audio (right). SoundMind therefore offers complete logical reasoning tasks with fully aligned content and response across text and audio modalities, supporting multimodal training and evaluation of large audio–language models.
@@
Large language models (LLMs) have made remarkable strides in reasoning capabilities through innovations such as Chain-of-Thought (CoT) prompting and specialized reasoning architectures. Recent models such as OpenAI’s o1 model (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025) have demonstrated exceptional performance on complex logical tasks (Kimi Team et al., 2025; Zhao et al., 2024; Yuan et al., 2025; Zhang et al., 2024, 2025a, 2025b; Han et al., 2024). Particularly noteworthy is Deepseek-R1’s rule-based reinforcement learning approach, which enables emergent reasoning without relying on traditional frameworks like Monte Carlo Tree Search (Wan et al., 2024) or process reward models (Lightman et al., 2023). This general reasoning paradigm has been successfully extended to the visual domain, where frameworks such as Visual-CoT (Shao et al., 2024) have significantly improved multimodal models’ cognitive abilities for image and video reasoning.
@@
Despite these advances in text and visual reasoning, there is a significant gap in audio reasoning and generation capabilities. While large audio-language models (LALMs) like Audio Flamingo (Kong et al., 2024) and Qwen2.5-Omni-7B (Xu et al., 2025) have made progress in audio understanding, end-to-end audio reasoning remains underdeveloped. This limitation stems primarily from two factors: (1) the simplicity of existing audio reasoning datasets (Suzgun et al., 2023; Kong et al., 2024), which often contain only brief textual labels without proper audio modality annotations, and (2) the technical challenges in maintaining reasoning coherence during long-duration audio generation. Current CoT methods applied to audio often lead to hallucinations and performance degradation when generating extended reasoning sequences. The lack of aligned audio-text annotations further impedes research on reasoning-driven audio generation, creating a bottleneck in developing LALMs with sophisticated reasoning capabilities.
@@
To address these challenges in audio reasoning, we introduce a comprehensive approach with two key components. First, we introduce SoundMind, a benchmark dataset specifically designed for the audio logical reasoning (ALR) task. It contains 6,446 audio–text aligned samples, each containing user content, step-by-step chain-of-thought reasoning, final answers, and the corresponding input–output speech audio (Figure 1). Built upon the LogiQA 2.0–NLI dataset (Liu et al., 2023), SoundMind preserves the full logical structure and augments it with parallel text–audio annotations, enabling models to learn reasoning grounded in audio. Second, we propose SoundMind-RL, a rule-based reinforcement learning algorithm that addresses the challenges of long-form audio reasoning generation. Drawing inspiration from the Logic-RL framework (Xie et al., 2025a), SoundMind-RL incorporates a strict format-based reward mechanism to prevent shortcut reasoning biased toward the textual modality. It leverages the REINFORCE++ algorithm (Hu et al., 2025) alongside the reward design principles from Deepseek-R1 for effective post-training. Our main contributions are summarized as follows:
@@
  • SoundMind Dataset: We release a high-quality audio logical reasoning dataset comprising 6,446 samples with user content, CoT reasoning, answers and corresponding audio input/output. It provides deep reasoning annotations in both text and audio modalities, serving as a valuable resource for RL-based training and evaluation of audio–language models.
@@
  • SoundMind-RL Algorithm: Taking into account the unique challenges of audio reasoning generation, we design a strict format-based reward mechanism that maintains reasoning coherence across modalities. By combining the improved REINFORCE++ algorithm with reward design principles from DeepSeek-R1, we fine-tune Qwen2.5-Omni-7B through reinforcement learning, achieving state-of-the-art results across three input–output modality configurations on the SoundMind benchmark.
@@
SoundMind (Ours) Speech 6446 1K
Table 1: Comparison of publicly available datasets for audio reasoning. Existing resources differ in scope: some provide speech data without transcripts, while others include transcripts but lack step-by-step reasoning annotations aligned with audio. These gaps make systematic study of multimodal reasoning challenging. SoundMind combines transcripts, chain-of-thought (CoT) annotations, and speech audio into a unified benchmark, offering a balanced resource for training and evaluating audio–language models.
2 Related Work
Open-Source Datasets for Audio Reasoning.
Research on audio reasoning remains limited by the scarcity of publicly available benchmarks, making systematic evaluation and comparison difficult. Existing corpora typically treat audio as descriptive input paired with text, yet their supervision remains purely textual, providing final annotations exclusively in text form (Suzgun et al., 2023; Kong et al., 2024; Ghosh et al., 2025). More recently, a number of studies have begun to recognize the critical role of chain-of-thought (CoT) prompting in audio reasoning datasets (Xie et al., 2025b), highlighting the importance of structured reasoning traces. To the best of our knowledge, SoundMind is among the first open-source resources where audio itself serves as the annotated modality for reasoning, providing parallel text–audio representations and enabling models to learn from and be evaluated on spoken reasoning traces.
@@ 2 Related Work
LLMs enhance their reasoning ability through in-context learning (ICL), which processes prompts within the surrounding context. CoT techniques further reinforce this capability. Prominent CoT approaches include Tree-of-Thought (ToT) (Yao et al., 2023), manually crafted few-shot CoT (Wei et al., 2022), and various automatic generation strategies (Jin et al., 2024). Recent works have also examined the necessity, theoretical underpinnings, and task-specific effectiveness of CoT reasoning (Sprague et al., 2025). OpenAI’s ol model (Jaech et al., 2024) has reignited interest in CoT prompting and is often paired with reinforcement learning-based training approaches (Hu et al., 2025; Xie et al., 2025a).
@@
Although current audio-language models (ALMs) have made progress in comprehension and real-time response, their capability in CoT-style reasoning remains underexplored. Our study addresses this research gap by systematically investigating the application of CoT techniques within ALMs.
3 Dataset Design and Construction
3.1 Key Characteristics
We introduce SoundMind, a unified benchmark for the audio logical reasoning (ALR) task, specifically developed to advance research on reasoning grounded in auditory input. In contrast to most existing datasets that provide only textual supervision, SoundMind offers both text-level and audio-level chain-of-thought (CoT) annotations, allowing models to learn and reason directly from spoken prompts and responses. The dataset emphasizes speech-based scenarios, making it well suited for applications such as dialogue understanding and auditory commonsense reasoning.
As shown in Table 1, existing audio datasets such as CoTA (Xie et al., 2025b), AudioSkills (Ghosh et al., 2025), LongAudio (Ghosh et al., 2025), and Big Bench Audio (Suzgun et al., 2023) provide valuable resources but remain limited in their support for auditory reasoning. Some lack audio-level annotations entirely, while others include audio but omit chain-of-thought (CoT) reasoning traces or detailed step-by-step rationales. For instance, although CoTA offers CoT-style annotations, it does not include audio transcripts, which limits its utility for training and evaluating models that operate directly on speech input. Likewise, AudioSkills and LongAudio supply a large quantity of audio data but no reasoning supervision, making them less suitable for investigating step-by-step inference or developing models that require explicit alignment between input and reasoning output.
Figure 2: Three-step pipeline for constructing SoundMind samples. Step 1: Structured logical triplets are converted into natural, conversational prompts through a Colloquialization module. Step 2: A large language model generates detailed chain-of-thought reasoning and final answers, providing rich supervision for reasoning tasks. Step 3: Both the user content and the generated reasoning–answer pairs are converted into speech using a text-to-speech (TTS) model, yielding two carefully aligned audio segments that together capture the full reasoning interaction.
The SoundMind dataset is constructed through a structured pipeline that transforms textual logical reasoning tasks into natural spoken audio interactions, thereby enabling the development of audio-based reasoning models. The complete pipeline is illustrated in Figure 2.
Although SoundMind contains fewer samples (6,446) than large-scale datasets such as AudioSkills, it is specifically designed for logic-oriented reasoning and provides over 1,074 hours of carefully annotated speech audio with both text-level and audio-level CoT supervision. This comprehensive annotation makes SoundMind a valuable benchmark for training and evaluating models that require structured, multimodal reasoning grounded in audio perception.
@@ 3.2 Data Generation Pipeline
Our pipeline takes as input structured logical triplets consisting of a major premise, a minor premise, and a conclusion. These are first processed by a Colloquialization module that rewrites the formal statements into natural, conversational prompts (e.g., “The major premise is …”, “Let’s figure out the logical connection …”). This step enhances the fluency of the resulting audio and more closely reflects realistic user queries.
@@
Table 2: Instructional prompts used at different positions during chain-of-thought generation. The system prompt specifies the overall task, output format, and encourages a careful, colloquial reasoning style suitable for speech synthesis. Additional contextual prompts placed before the major premise and after the conclusion provide further guidance, helping the model interpret logical relations and produce well-structured reasoning traces.
Figure 3: An illustrative example from the SoundMind dataset. Each sample contains three components: (1) User Content, which includes a natural-language prompt and a structured logical triplet (major premise, minor premise, conclusion); (2) Chain-of-Thought, a step-by-step reasoning trace generated by a large language model, explaining the entailment decision; and (3) the final Answer. All components are provided in both text form and synthesized speech audio, enabling fully multimodal training and evaluation of audio–language models.
The colloquialized content and instructions are merged into a single User Content block and synthesized into speech audio using the high-fidelity text-to-speech model MegaTTS 3 (Jiang et al., 2025), forming the input side of the dataset. The prompts used in this process are listed in Table 2.
To generate the output, we employ the large language model DeepSeek-R1 to produce step-by-step CoT reasoning and final answers, providing richer supervision for model training. The generated reasoning and answers are then synthesized into speech audio with MegaTTS 3 (Jiang et al., 2025), yielding the corresponding spoken response.
As a result, each SoundMind instance consists of two carefully aligned audio segments: one representing the user query and the other presenting the detailed step-by-step reasoning and the corresponding final answer, as illustrated in Figure 3.
4 SoundMind-RL Algorithm
@@
To ensure the correctness of answer formatting, we first require that the token “Answer:” must appear within the last five characters of the model response. Considering that the SoundMind dataset covers both text and audio modalities, we design two specific format scoring methods (denoted as $S_{\text{format}}$), calculated as follows:
@@
We additionally introduce a reward evaluation based on the reasoning length of the model response. This score ($S_{\text{len}}$) is computed by comparing the ratio of the model output length to the reference reasoning length. Given that the SoundMind dataset provides supervision in both text and audio modalities, we design two distinct length evaluation methods, defined as follows:
@@
where $L$ denotes the length of text tokens, and $T$ denotes the length of audio tokens.
@@
In our setting, the policy $\pi_{\theta}$ corresponds to a large-scale ALM, which receives an audio question as input and produces a reasoning response in either text or audio form. To optimize the model with the above composite reward, we adopt the REINFORCE++ Hu et al. (2025)—a clipped policy-gradient method that eliminates the need for a value (critic) network while leveraging PPO-style stability and sample efficiency. Specifically, REINFORCE++ updates the policy by maximizing the following objective:
@@ 5 Experiments
Training 2326 2858 182 1683 62.33 608.42
Test 296 360 158 1424 57.90 586.51
Validation 264 342 155 1426 57.85 586.39
Table 3: Detailed statistics of the SoundMind dataset across training, validation, and test splits. “# Entailed” and “# Not-ent.” denote the number of samples labeled as entailed and not-entailed. “Avg. Inp./Out. Tok.” reports the mean token counts for the input and outputs, respectively. “Avg. Inp./Out. Dur. (s)” provides the average duration (in seconds) of the input audio and the corresponding output audio. Together, these statistics highlight the dataset’s scale, near-balanced class distribution, and focus on long-form reasoning supervision.
@@
Figure 4: Distribution of audio durations across dataset splits and entailment labels. Subfigures (a–c) show histograms for “Entailed” samples in the training, test, and validation sets, while (d–f) show the corresponding distributions for “Not-entailed” samples. These plots illustrate the broad coverage of audio durations and the consistent duration profiles across splits, highlighting that the SoundMind dataset offers both diversity and balance for evaluating model performance across different temporal contexts.